OpenStudy (anonymous):

10.02] Which ratio forms a proportion with 6 over 15.? 8 over 10. 5 over 2. 3 over 12. 8 over 20.

OpenStudy (anonymous):

\[\frac{ 2 }{ 5 }=\frac{ 4 }{ 10 }=\frac{ 6 }{ 15 }=\frac{ 8 }{ 20 }=\frac{ 10 }{ 25 }\]

OpenStudy (anonymous):

So, the answer is?

OpenStudy (anonymous):

2/5, but that isn't there

OpenStudy (anonymous):

i didnt say 2/5 was the answer; answer is hidden go get it

OpenStudy (anonymous):

Oh, wait , sorry! I got it!